A discrete random variable X is believe to follow a binomial distribution b(x; 5, p). If P(X = 0) = 243/1024, then find P(X = 3).

1
Expert's answer
2022-05-23T23:48:47-0400
"P(X=0)=\dbinom{5}{0}(p)^{0}(1-p)^{5-0}=(1-p)^{5}"

"=\dfrac{243}{1024}"

"1-p=\dfrac{3}{4}"

"p=\dfrac{1}{4}"

"P(X=3)=\dbinom{5}{3}(\dfrac{1}{4})^{3}(\dfrac{3}{4})^{5-3}=\dfrac{45}{512}"